# Out-of-equlibrium inference of stochastic model parameters through population data from generic consumer-resource dynamics

### v1

@@ II Multi-resource Holling Type II feeding dynamics
The first channel represents feeding interactions, by taking into account different resource preferences determined by distinct attack rates ($\alpha_{i})$ for the resources. The second reaction models consumer-resource compound relaxation, which occurs at rates $n_{i}$ (depending on resource type). Mean handling times, given by $\nu_{i}^{-1}$, are also characteristic for individual consumers feeding on each resource type. Our aim in this section is to solve analytically the stochastic dynamics, by obtaining the joint probability distribution of the number of searching predators, as well as the number of consumers handling on each resource, both out-of-equilibrium and at stationarity.
@@
The master equation determines the temporal evolution of the probability of finding a given discrete configuration of consumer types, $(n_{A},n_{1},\dots,n_{S})$, at a given time $t$. As in [27], we assume chemostatic conditions, so the overall number of consumers, $n_{A}^{0}$, is kept constant. Therefore, the number of predators $n_{A}$ is immediately determined by the number of consumers handling on each resource, $\{n_{i}\}$, via Eq. (4). Accordingly, we define the state vector of the stochastic process as $\bm{n}=(n_{1},\dots,n_{S})$, considering only variations in handling predator numbers. Given a configuration vector $\bm{n}$, the number of free predators is determined as $n_{A}=n_{A}^{0}-\sum_{i=1}^{S}n_{i}$.
@@
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\alpha_{i}\frac{n_{A}n_{R_{i}}}{N},$ (5)
$T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$
for $i=1,\dots,S$, each line corresponding to chemical reactions (2) and (3), respectively. They account for the conditional probability for transition $\bm{n}\to\bm{n}+\bm{e}_{i}$, i.e, the event of a consumer that starts handling on resource $i$, as well as the conditional probability for transition $\bm{n}\to\bm{n}-\bm{e}_{i}$, which represents the relaxation of the compound formed by the consumer and the $i$-th resource. For handling predators formation, we assume also chemostatic conditions for resources levels, meaning that resource densities $x_{R_{i}}:=n_{R_{i}}/N$ are kept constant —here $N$ stands for the number of sites of the system (system volume). In addition, a law of mass action is assumed because the probability of consumer-resource encounter has to be proportional to resource availability. The effective per capita attack rate of a free consumer at a given density level of the $i$-th resource type is a resource-specific constant parameter defined by $\theta_{i}:=\alpha_{i}x_{R_{i}}$. The second transition probability represents $AR_{i}$ compounds relaxation, leading to free predators after a mean handling time $\nu_{i}^{-1}$. This probability is proportional to the number of handling predators $n_{i}$ if one of the $AR_{i}$ compounds stops handling the resource.
@@
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\theta_{i}\bigg{(}n_{A}^{0}-\sum_{j=1}^{S}n_{j}\bigg{)},$ (6)
$T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$
@@
$\frac{dP_{\bm{n}}}{dt}$ $=\sum_{i=1}^{S}\left\{T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}+T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}\right.$ (7)
$-\left.\left[T(\bm{n}+\bm{e_{i}}|\bm{n})+T(\bm{n}-\bm{e_{i}}|\bm{n})\right]P_{\bm{n}}\right\}.$
@@
II.1 Steady-state distribution
@@
$n_{A}$ being the number of searching consumers [28]. One can take advantage of knowing that distribution to rigorously derive the functional response associated to the single reaource case ($S=1$), which turns out to be the Holling type II functional response [28]. We here extend the binomial distribution result to a multi-resource case by solving the master equation for stationarity, which yields an analytical expression for the joint steady-state distribution $P_{\bm{n}}$. We can do this because transition probability rates (6) are linear in population numbers [29, 30]. As a result, the single-resource binomial distribution is extended to a multi-resource case, leading to a multinomial distribution for the joint probability distribution of observing a vector $(n_{A},\bm{n})$ of the populations of behavioral types, as we show next.
@@
$P_{\bm{n}}=\frac{\nu_{i}}{\theta_{i}}\frac{n_{i}+1}{n_{A}^{0}-\sum_{k=1}^{S}n_{k}}P_{\bm{n}+\bm{e}_{i}},$ (12)
$P_{\bm{n}}=\frac{\nu_{j}}{\theta_{j}}\frac{n_{j}+1}{n_{A}^{0}-\sum_{k=1}^{S}n_{k}}P_{\bm{n}+\bm{e}_{j}},$
@@
$P_{\bm{n}}$ $=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}}{n_{1}}$ (18)
×$\times\binom{n_{1}+n_{2}+n_{3}}{n_{1}+n_{2}}\times P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$
@@
$P_{\bm{n}}$ $=\left(\frac{\theta_{1}}{\nu_{1}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\right)^{n_{2}}\left(\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}$ (19)
×$\times\binom{n_{1}+n_{2}+n_{3}}{n_{1},n_{2},n_{3}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$
where we have used known relations between the binomial and the multinomial coefficients, where the latter are defined by $\binom{m}{s_{1},\dots,s_{q}}=\frac{m!}{\prod_{k=1}^{q}s_{k}!}$ (with $s_{1}+\dots+s_{q}=m$). Here, probabilities are expressed in terms of configurations in which no individual consumers feed on resources $1$ and $2$.
@@
$P_{\bm{n}}$ $=\binom{\sum_{k=1}^{S}n_{k}}{n_{1},\dots,n_{S}}\left(\frac{\nu_{S}}{\theta_{S}}\right)^{\sum_{k=1}^{S-1}n_{k}}\left[\prod_{k=1}^{S-1}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]$ (20)
×$\times P_{(0,\dots,0,\sum_{k=1}^{S}n_{k})}.$
@@
According to the multinomial theorem [31], the sum above can be explicitly expressed as
@@
The joint distribution of the vector of behavioral types population abundances, $(n_{A},n_{1},\dots,n_{S})$, is given by Eqs. (24) and (31). It is a multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$, as defined in Eq. (26), with $n_{A}^{0}$ trials and probabilities given by
$p_{i}=\frac{\theta_{i}/\nu_{i}}{1+\sum_{k=1}^{S}\theta_{k}/\nu_{k}}$ (32)
@@
This automatically yields the multi-resource Holling type II functional response extension (see [28] for the derivation using a different methodology).
In addition to the joint, steady-state probability distribution of different behavioral types, by summing over the corresponding variables we can obtain the marginal distributions of each single type, which are binomials $B(n_{A}^{0},p_{i})$ [26]. In particular, from our derivation above we find that the distribution of the total number of handling consumers is a binomial $B(n_{A}^{0},1-p_{A})$: using Eq. (31) and (33), the probability that $\sum_{k=1}^{S}n_{k}=q$, which is given by Eq. (30), can be rewritten as
@@
as expected [28].
II.2 Out-of-equilibrium probability distribution
Calculating the out-of-equilibrium probability distribution involves integrating the master equation (7) as function of time. Usually this equation has to be integrated numerically. In this case, however, an analytical solution is feasible because transition probability rates depend linearly on population numbers [29, 30]. We solve the master equation by introducing a generating function, which is basically a multi-dimensional zeta transform of the probability distribution $P_{\bm{n}}(t)$. This transformation converts the infinite system of coupled differential equations into a partial differential equation.
The generating function $G(\bm{z},t)$ is defined as
where the vector $\bm{z}=(z_{1},\dots,z_{S})$ is conjugate to the state vector $\bm{n}=(n_{1},\dots,n_{S})$. This vector $\bm{z}$ can take values over $\mathbb{R}^{S}$. Then, by multiplying the master equation, Eq. (7), by the product $z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}$, and summing over the population variables $n_{i}$ ($i=1,\dots,S$) over non-negative integers, we obtain the following PDE for the generating function $G$,
@@ II.2 Out-of-equilibrium probability distribution
We leave the details of the derivation of Eq. (40) from (7) to Appendix A.
@@
i.e., initially the state of the system is defined by the population vector $\bm{\hat{n}}=(\hat{n}_{1},\dots,\hat{n}_{S})$, where $\hat{n}_{i}$ stands for the initial number of predators handling resource $i$, and there are $\hat{n}_{A}=n_{A}^{0}-\sum_{j=1}^{S}\hat{n}_{j}$ free predators —we assume that $n_{A}^{0}-\sum_{j=1}^{S}\hat{n}_{j}\geq 0$ holds. In the initial condition above, $\delta_{n,m}$ stands for the Kronecker symbol, i.e., $\delta_{n,m}=1$ if $n=m$ and 00 otherwise. By definition, the initial condition reduces to
@@
Now we solve the first-order PDE (40) using the method of characteristics curves. An equivalent way of writing the initial PDE in a very compact form is as the following dot product, $\bm{v}\cdot\bm{\nabla}H=0$, where the solution $g=G(\bm{z},t)$ is viewed in the $(S+2)$-dimensional space $(\bm{z},t,g)$ as the manifold $H(\bm{z},t,g)=G(\bm{z},t)-g=0$, and vector $\bm{v}$ is defined as
@@
Here the row vector $\bm{w}=(w_{i})$ has $S$ components defined by $w_{i}=\nu_{i}(1-z_{i})+m(\bm{z})z_{i}$, for $i=1,\dots,S$. Finding the integral surface $H$ amounts to solving the PDE.
@@
In the particular case in which all the relaxation rates are equal, $\nu_{1}=\dots=\nu_{S}=:\nu$, the system decouples because the second equation is closed for $m$, according to Eq. (41), so we can find the time dependence of function $m(t):=m(\bm{z}(t))$. This amounts to assuming that the predator spends the same average time $\nu^{-1}$ handling any type of resource. Therefore, the equation for $m(t)$ reduces to a logistic growth equation if the assumption of equal handling times (equal relaxation rates) holds:
@@
for $K$ an arbitrary constant to be calculated as function of the initial conditions, $\hat{z}_{j}:=z_{j}(0)$. Then solving for $K$ in Eq. (50) for $t=0$ gives
@@
where we have defined $R_{0}:=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$. Observe that the initial condition vector $\bm{\hat{z}}$ depends on the vector of $z$-variables, $\bm{z}$, via the explicit solution of the differential equations for $z_{i}$ in Eq. (46). Because the solution of the system of ODEs satisfied by the characteristic curves will depend on these initial conditions $\bm{\hat{z}}$, we have to substitute $\hat{z}_{j}$ as function of $\bm{z}$ at the end of the calculation in order to get the desired expression for the generating function $G(\bm{z},t)$.
@@
Recall that the desired dependence on $\bm{z}$ is obtained once we make the substitution $\hat{z}_{j}=\hat{z}_{j}(\bm{z},t)$. So it only remains to find the dependence of the initial conditions $\hat{z}_{j}$ in terms of $z$-variables.
@@
where we have defined $R=R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$. Because $\hat{z}_{j}$ are at the end functions of $\bm{z}$, $K$ becomes itself a function of $\bm{z}$, which can be determined as follows. From Eq. (51) we can eliminate $R_{0}=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$ in terms of $K$ as $R_{0}=(KQ-\nu)/(1+K)$ which, upon substitution into the last expression, yields
@@
Finally, an explicit expression of initial conditions $\hat{z}_{j}$ as function of $\bm{z}$ is determined by substituting $K(\bm{z})$, given by Eq. (58), into (55), which results in an equation from which we can eliminate $\hat{z}_{j}$ in terms of $\bm{z}$. This process gives
@@
$G(\bm{z},t)=\left[\frac{R({\bm{z}})+\nu+(Q-R({\bm{z}}))e^{-(\nu+Q)t}}{\nu+Q}\right]^{n_{A}^{0}}\\ \times\prod_{j=1}^{S}\left[1-\left(1+\frac{\nu}{Q}\right)\frac{Q-R(\bm{z})+(R(\bm{z})-Qz_{j})e^{Qt}}{Q-R(\bm{z})+(R(\bm{z})+\nu)e^{(\nu+Q)t}}\right]^{\hat{n}_{j}}.$ (60)
We recall here the definitions $R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$ and $Q:=\sum_{j=1}^{S}\theta_{j}$, and that the condition $n_{A}^{0}-\sum_{j=0}^{S}\hat{n}_{j}\geq 0$ must hold for the last expression to make sense. This function satisfies the initial condition (43) as well as the normalization condition $G(\bm{1},t)=1$. The generating function has to be expanded in power series of $z_{i}$ in order to obtain the explicit expression of the probability distribution $P_{\bm{n}}(t)$. Such task is hard in general.
@@
for $j=1,\dots,S$. It is easy to check that $0\leq p_{j}(t)\leq 1$ for $j=0,1,\dots,S$ and arbitrary time $t$. In addition, it holds that $\sum_{j=0}^{S}p_{j}(t)=1$ for any time. Therefore, the parameter combinations $p_{j}(t)$ can be interpreted as probabilities. As we show immediately below, they are precisely the probabilities associated to the $j$-th population in a multinomial distribution.
@@
$G(\bm{z},t)$ $=\biggl{(}p_{0}(t)+\sum_{j=1}^{S}p_{j}(t)z_{j}\biggr{)}^{n_{A}^{0}}$ (63)
$=\sum_{n_{A}=0}^{n_{A}^{0}}\binom{n_{A}^{0}}{n_{A}}p_{0}^{n_{A}}\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}.$
Then, by the multinomial theorem [31],
@@
This agrees with the results obtained at equilibrium, in the limit $t\to\infty$. As explained in the previous subsection, we also found a multinomial distribution in the stationary regime. However, it is not necessarily true that the out-of-equilibrium distribution has to be multinomial in any case, although it must asymptotically converge to a multinomial. This is obvious from Eq. (60): in general, the generating function is no longer the one associated to a multinomial distribution, which is given by Eq. (61), but it has the correct limit as $t\to\infty$. Indeed, if we take this limit on both sides of Eq. (60) we get
$\lim_{t\to\infty}G(\bm{z},t)=\left[\frac{\nu+R({\bm{z}})}{\nu+Q}\right]^{n_{A}^{0}}=\biggl{(}p_{A}+\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}},$ (66)

### v2

@@ 3 Multi-resource Holling Type II feeding dynamics
The first channel represents feeding interactions, by taking into account different resource preferences determined by distinct attack rates ($\alpha_{i})$ for the resources. The second reaction models consumer-resource compound relaxation, which occurs at rates $\nu_{i}$ (depending on resource type). Mean handling times, given by $\nu_{i}^{-1}$, are also characteristic for individual consumers feeding on each resource type. Our aim in this section is to solve analytically the stochastic model, by obtaining the joint probability distribution of the number of searching predators, as well as the number of consumers handling on each resource, both out-of-equilibrium and at stationarity.
@@
The master equation determines the temporal evolution of the probability of finding a given discrete configuration of consumer types, $(n_{A},n_{1},\dots,n_{S})$, at a given time $t$. As in Palamara et al. (2021), we assume chemostatic conditions, so the overall number of consumers, $n_{A}^{0}$, is kept constant. Therefore, the number of predators $n_{A}$ is immediately determined by the number of consumers handling on each resource, $\{n_{i}\}$, via Eq. (11). Accordingly, we define the state vector of the stochastic process as $\bm{n}=(n_{1},\dots,n_{S})$, considering only variations in handling predator numbers. Given a configuration vector $\bm{n}$, the number of free predators is determined as $n_{A}=n_{A}^{0}-\sum_{i=1}^{S}n_{i}$.
@@
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\alpha_{i}\frac{n_{A}n_{R_{i}}}{N},\qquad T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$ (12)
for $i=1,\dots,S$, each one corresponding to chemical reactions (9) and (10), respectively. They account for the conditional probability for transition $\bm{n}\to\bm{n}+\bm{e}_{i}$, i.e, the event of a consumer that starts handling on resource $i$, as well as the conditional probability for transition $\bm{n}\to\bm{n}-\bm{e}_{i}$, which represents the relaxation of the compound formed by the consumer and the $i$-th resource. For handling predators formation, we assume also chemostatic conditions for resources levels, meaning that resource densities $x_{R_{i}}:=n_{R_{i}}/N$ are kept constant —here $N$ stands for the number of sites of the system (system volume). In addition, a law of mass action is assumed because the probability of consumer-resource encounter has to be proportional to resource availability. The effective per capita attack rate of a free consumer at a given density level of the $i$-th resource type is a resource-specific constant parameter defined by $\theta_{i}:=\alpha_{i}x_{R_{i}}$. The second transition probability represents $AR_{i}$ compounds relaxation, leading to free predators after a mean handling time $\nu_{i}^{-1}$. This probability is proportional to the number of handling predators $n_{i}$ if one of the $AR_{i}$ compounds stops handling the resource.
@@
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\theta_{i}\bigg{(}n_{A}^{0}-\sum_{j=1}^{S}n_{j}\bigg{)},\qquad T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$ (13)
@@
$\frac{dP_{\bm{n}}}{dt}=\sum_{i=1}^{S}\left\{T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}+T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}-\left[T(\bm{n}+\bm{e_{i}}|\bm{n})+T(\bm{n}-\bm{e_{i}}|\bm{n})\right]P_{\bm{n}}\right\}.$ (14)
@@
Our aim is to estimate properly feeding rates $\{\alpha_{1},\dots,\alpha_{S},\nu_{1},\dots,\nu_{S}\}$. We first compute the steady-state distribution for arbitrary parameter values, and show that the distribution of feeding types is multinomial. To do that, we use a detailed-balance condition satisfied by transition probabilities. Second, using the generating function formalism, we find an analytical solution for out-of-equilibrium probabilities $P_{\bm{n}}(t)$ for the particular case of equal relaxation rates, $\nu_{1}=\dots=\nu_{S}$, and for an initial state formed only by searching predators. In the second case, we also show that the distribution is multinomial. From the form of the distributions (and the generating function), it is apparent that the model is perfectly identifiable out of equilibrium.
3.1 Steady-state distribution
@@
$n_{A}$ being the number of searching consumers (Palamara et al., 2022). One can take advantage of knowing that distribution to rigorously derive the functional response associated to the single reaource case ($S=1$), which turns out to be the Holling type II functional response (Palamara et al., 2022). We here extend the binomial distribution result to a multi-resource case by solving the master equation at stationarity, which yields an analytical expression for the joint steady-state distribution $P_{\bm{n}}$. We can do this because transition probability rates (13) are linear in population numbers (Kendall, 1949; Nisbet and Gurney, 2004). As a result, the single-resource binomial distribution is extended to a multi-resource case, leading to a multinomial distribution for the joint probability distribution of observing a vector $(n_{A},\bm{n})$ of the populations of behavioral types, as we show next.
@@
$P_{\bm{n}}=\frac{\nu_{i}}{\theta_{i}}\frac{n_{i}+1}{n_{A}^{0}-\sum_{k=1}^{S}n_{k}}P_{\bm{n}+\bm{e}_{i}},\qquad P_{\bm{n}}=\frac{\nu_{j}}{\theta_{j}}\frac{n_{j}+1}{n_{A}^{0}-\sum_{k=1}^{S}n_{k}}P_{\bm{n}+\bm{e}_{j}},$ (19)
@@
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}}{n_{1}}\binom{n_{1}+n_{2}+n_{3}}{n_{1}+n_{2}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$ (25)
@@
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\right)^{n_{2}}\left(\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}+n_{3}}{n_{1},n_{2},n_{3}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$ (26)
where we have used known relations between the binomial and the multinomial coefficients, the latter being defined as $\binom{m}{s_{1},\dots,s_{q}}=\frac{m!}{\prod_{k=1}^{q}s_{k}!}$ (with $s_{1}+\dots+s_{q}=m$). Here, probabilities are expressed in terms of configurations in which no individual consumers feed on resources $1$ and $2$.
@@
$P_{\bm{n}}=\binom{\sum_{k=1}^{S}n_{k}}{n_{1},\dots,n_{S}}\left(\frac{\nu_{S}}{\theta_{S}}\right)^{\sum_{k=1}^{S-1}n_{k}}\left[\prod_{k=1}^{S-1}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]P_{(0,\dots,0,\sum_{k=1}^{S}n_{k})}.$ (27)
@@
According to the multinomial theorem (Knuth, 1997), the sum above can be explicitly expressed as
@@
The joint distribution of the vector of population abundances of different behavioral types, $(n_{A},n_{1},\dots,n_{S})$, is given by Eqs. (31) and (38). It is a multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$, as defined in Eq. (33), with $n_{A}^{0}$ trials and probabilities given by
$p_{i}=\frac{\theta_{i}}{\nu_{i}}\frac{1}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}}$ (39)
@@
This automatically yields the multi-resource Holling type II functional response extension (see Palamara et al. (2022) for the derivation using a different methodology).
In addition to the joint, steady-state probability distribution of different behavioral types, by summing over the corresponding variables we can obtain the marginal distributions of each single type, which are binomials $B(n_{A}^{0},p_{i})$ (Loève and Loève, 1977). In particular, from our derivation above we find that the distribution of the total number of handling consumers is a binomial $B(n_{A}^{0},1-p_{A})$: using Eq. (38) and (40), the probability that $\sum_{k=1}^{S}n_{k}=q$, which is given by Eq. (37), can be rewritten as
@@
as expected (Palamara et al., 2022).
If we were to use the multinomial steady-state distribution (33) with probabilities given by Eqs. (39) and (40), to estimate feeding rates vía ML (as above, we assume that $n_{A}^{0}$ is fixed and not subject to MLE), we could write the following exhaustive summary: $\bm{\kappa}=\left(\theta_{1}/\nu_{1},\dots,\theta_{S}/\nu_{S}\right)$, which is tantamount to $\bm{\kappa}=\left(\alpha_{1}/\nu_{1},\dots,\alpha_{S}/\nu_{S}\right)$ because resource densities $x_{R_{i}}$ are kept constant (chemostatic conditions). Since the log-likelihood is determined by $S$ ratios and there are $2S$ individual rates ($\alpha_{1},\dots,\alpha_{S},\nu_{1},\dots,\nu_{S}$) to be estimated, the model is structurally unidentifiable (parameter redundant) if MLE is conducted using the steady-state distribution.
3.2 Out-of-equilibrium probability distribution
Calculating the out-of-equilibrium probability distribution involves integrating the master equation (14) as a function of time. Usually this equation has to be integrated numerically. In this case, however, an analytical solution is feasible because transition probability rates depend linearly on population numbers (Kendall, 1949; Nisbet and Gurney, 2004). We solve the master equation by introducing the generating function defined in Eq. (2). Transforming $P_{\bm{n}}(t)$ into the generating function $G(\bm{z},t)$ converts the infinite system of coupled differential equations into a partial differential equation.
We start by obtaining the PDE satisfied by the generating function $G(\bm{z},t)$ defined in Eq. (2). By multiplying the master equation, Eq. (14), by the product $z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}$, and summing over the population variables $n_{i}$ ($i=1,\dots,S$) over non-negative integers, we obtain the following PDE for the generating function $G$,
@@ 3.2 Out-of-equilibrium probability distribution
We leave the details of the derivation of Eq. (46) from (14) to A.
@@
i.e., initially the state of the system is defined by the population vector $\bm{\hat{n}}=(\hat{n}_{1},\dots,\hat{n}_{S})$, where $\hat{n}_{i}$ stands for the initial number of predators handling resource $i$, and there are $\hat{n}_{A}=n_{A}^{0}-\sum_{j=1}^{S}\hat{n}_{j}$ free predators —we assume that $n_{A}^{0}-\sum_{j=1}^{S}\hat{n}_{j}\geq 0$ holds. In the initial condition above, $\delta_{n,m}$ stands for the Kronecker symbol, i.e., $\delta_{n,m}=1$ if $n=m$ and 00 otherwise. In terms of $G(\bm{z},t)$, the initial condition reduces to
@@
Now we solve the first-order PDE (46) using the method of characteristics curves. An equivalent way of writing the initial PDE in a very compact form is as the following dot product, $\bm{v}\cdot\bm{\nabla}H=0$, where the solution $g=G(\bm{z},t)$ is viewed in the $(S+2)$-dimensional space $(\bm{z},t,g)$ as the manifold $H(\bm{z},t,g)=G(\bm{z},t)-g=0$, and the vector field $\bm{v}$ is defined as
@@
Here the row vector $\bm{w}=(w_{i})$ has $S$ components defined as $w_{i}=\nu_{i}(1-z_{i})+m(\bm{z})z_{i}$, for $i=1,\dots,S$. Finding the integral surface $H$ amounts to solving the PDE.
@@
In the particular case in which all the relaxation rates are equal, $\nu_{1}=\dots=\nu_{S}=:\nu$, the system decouples because the second equation is closed for $m$, according to Eq. (47), so we can find the time dependence of function $m(t):=m(\bm{z}(t))$. The condition of equal relaxation rates amounts to assuming that the predator spends the same average time $\nu^{-1}$ handling any type of resource. Therefore, the equation for $m(t)$ reduces to a logistic growth equation if the assumption of equal handling times (equal relaxation rates) holds:
@@
for $K$ an arbitrary constant to be calculated as function of the initial conditions, $\hat{z}_{j}:=z_{j}(0)$. Then solving for $K$ in Eq. (56) at $t=0$ gives
@@
where we have defined $R_{0}:=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$. Observe that the initial condition vector $\bm{\hat{z}}$ depends on the vector of $z$-variables, $\bm{z}$, via the explicit solution of the differential equations for $z_{i}$ in Eq. (52). Because the solution of the system of ODEs satisfied by the characteristic curves will depend on the vector $\bm{\hat{z}}$ of initial conditions, we have to substitute $\hat{z}_{j}$ as function of $\bm{z}$ at the end of the calculation in order to get the desired expression for the generating function $G(\bm{z},t)$.
@@
Recall that the desired dependence on $\bm{z}$ is obtained once we make the substitution $\hat{z}_{j}=\hat{z}_{j}(\bm{z},t)$. So it only remains to find the dependence of the initial conditions $\hat{z}_{j}$ in terms of $\bm{z}$.
@@
where we have defined $R=R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$. Because $\hat{z}_{j}$ are at the end functions of $\bm{z}$, $K$ becomes itself a function of $\bm{z}$, which can be determined as follows. From Eq. (57) we can solve for $R_{0}=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$ in terms of $K$ as $R_{0}=(KQ-\nu)/(1+K)$ which, upon substitution into the last expression, yields
@@
Finally, an explicit expression of initial conditions $\hat{z}_{j}$ as function of $\bm{z}$ is determined by substituting $K(\bm{z})$, given by Eq. (64), into (61), which results in an equation from which we can eliminate $\hat{z}_{j}$ in terms of $\bm{z}$. This manipulation gives
@@
$G(\bm{z},t)=\left[\frac{R({\bm{z}})+\nu+(Q-R({\bm{z}}))e^{-(\nu+Q)t}}{\nu+Q}\right]^{n_{A}^{0}}\prod_{j=1}^{S}\left[1-\left(1+\frac{\nu}{Q}\right)\frac{Q-R(\bm{z})+(R(\bm{z})-Qz_{j})e^{Qt}}{Q-R(\bm{z})+(R(\bm{z})+\nu)e^{(\nu+Q)t}}\right]^{\hat{n}_{j}}.$ (66)
Recall the definitions $R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$ and $Q:=\sum_{j=1}^{S}\theta_{j}$, and that the condition $n_{A}^{0}-\sum_{j=0}^{S}\hat{n}_{j}\geq 0$ must hold for the last expression to make sense. This function satisfies the initial condition (49) as well as the normalization condition $G(\bm{1},t)=1$. The generating function has to be expanded in power series of $z_{i}$ in order to obtain the explicit expression of the probability distribution $P_{\bm{n}}(t)$. Such task is hard in general. As shown by the derivation itself, it is very difficult to find the out-of-equilibrium distribution from the master equation, in contrast with the solution at stationarity: for the former, we cannot find it unless assuming equal handling rates, and we cannot carry out the series expansion to obtain the probability distribution $P_{\bm{n}}(t)$ for arbitrary initial conditions.
@@
for $j=1,\dots,S$. It is easy to check that $0\leq p_{j}(t)\leq 1$ for $j=0,1,\dots,S$ and arbitrary time $t$. In addition, it holds that $\sum_{j=0}^{S}p_{j}(t)=1$ for any time. Therefore, the parameter combinations $p_{j}(t)$ can be interpreted as probabilities. As we immediately show below, they are precisely the probabilities associated to the $j$-th population in a multinomial distribution.
@@
$G(\bm{z},t)=\biggl{(}p_{0}(t)+\sum_{j=1}^{S}p_{j}(t)z_{j}\biggr{)}^{n_{A}^{0}}=\sum_{n_{A}=0}^{n_{A}^{0}}\binom{n_{A}^{0}}{n_{A}}p_{0}^{n_{A}}\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}.$ (69)
Then, by the multinomial theorem (Knuth, 1997), it holds
@@
This result agrees with those obtained at equilibrium, in the limit $t\to\infty$. As explained in the previous subsection, we also found a multinomial distribution in the stationary regime. However, the out-of-equilibrium distribution is not multinomial in general: note that the generating function for non-zero initial populations, Eq. (66), does not match the one one associated to a multinomial distribution, which is given by Eq. (67), unless there are no handling predators initially. $P_{\bm{n}}(t)$ follows a multinomial distribution only if $\bm{\hat{n}}=\bm{0}$. Irrespective of the set of initial conditions, however, the generating function (66) has the correct limit as $t\to\infty$. Indeed, if we take this limit on both sides of Eq. (66) we get
$\lim_{t\to\infty}G(\bm{z},t)=\left(\frac{\nu+R({\bm{z}})}{\nu+Q}\right)^{n_{A}^{0}}=\biggl{(}p_{A}+\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}},$ (72)